At the beginning of 2011, Bonds Company had total assets of $650,000 and total liabilities of $390,000. Answer each of the following questions.                              
1. If total assets increased $60,000 and stockholders' equity decreased $90,000 during the year, determine the amount of total liabilities at the end of the year.                              
2. During the year, total liabilities decreased $75,000 and stockholders' equity increased $50,000. Compute the amount of total assets at the end of the year.                              
3. If total assets decreased $100,000 and total liabilities increased $55,000 during the year, determine the amount of stockholders' equity at the end of the year.                              
Answer #1
Beginning stockholders' Equity 260000 =650000-390000
1
Total assets 710000 =650000+60000
Less: Stockholders' equity 170000 =260000-90000
Total liabilities at the end of the year 540000
2
Total liabilities 315000 =390000-75000
Add: Stockholders' equity 310000 =260000+50000
Total assets at the end of the year 625000
3
Total assets 550000 =650000-100000
Less: Total liabilities 445000 =390000+55000
Stockholders' equity at the end of the year 105000